# A globally convergent gradient method with momentum

### v2

@@ 6.2 Hessian estimation in subspace by interpolation
$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$
@@ 7 Computational experiments
In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work. Specifically, we considered to this aim a benchmark of 163 problems from the CUTEst test-suite [28]. In particular, we included in our test set all 183 problems in the CUTEst collection that are 1) unconstrained, 2) regular, 3) with a number of variables $n$ that is greater or equal than 1000 (except for the problems LUKSAN12LS, LUKSAN13LS, LUKSAN14LS having 998 variables) or that is user definable; in the latter case, we set the number of variables to 1000 or as close as possible to 1000. In Table 1 those 183 problems are reported along with their dimensions.
@@
The code for all the experiments described in this work was written in Python 3.9, only exploiting numpy and scipy libraries. The software is available at:
Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq. 28 holds. As for the parameters, we set $\delta=0.5,\;\gamma=10^{-5}$.
The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration. As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
As a baseline for comparisons, we took into account classical nonlinear conjugate gradient methods [29] and the L-BFGS algorithm [30], which can arguably be considered the state-of-the-art for the solution of smooth unconstrained nonlinear optimization problems. For both algorithms, we considered the very efficient implementations available through the scipy library.
The results of the experiments are shown in the form of performance profiles [31]. As performance metrics, we considered both the runtime and the number of iterations.
In the following, we report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6. More in detail, we denote by GMM1, GMM2, GMM3 the algorithm where $H_{k}$ is computed as described in subsection 6.1, 6.2 and 6.3, respectively.
Figure 1: Performance profiles on all 163 problems.
@@
In fig. 1, we report the performance profiles in terms of time and number of iterations for the comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 163 problems. As we can see, the best performing solver is GMM2 both in terms of time and iterations. We can also say that GMM3 is the second best solver in terms of time even though it is the worst one in terms of iterations. Both solvers GMM2 and GMM3 are better than CGscipy in terms of time. In doing these considerations, we must also take into account that it might happen that solvers converge toward different points. Indeed, in table 2 we report the number of wins, i.e., the number of problems where each solver finds the best function value.
Table 2: Number of wins, i.e., the number of problems where each solver finds the best function value, on all 163 test problems. GMM variants and CGscipy solvers are considered.
Solver # wins
GMM1 71
GMM2 58
GMM3 26
CGscipy 16
@@
where $\cal S$ is the set of solvers. From fig. 2, we note that the performance of GMM1 becomes closer to that of the conjugate gradient method in terms of computational time, whereas GMM2 and GMM3 continue to be clearly superior.
@@
Figure 3: Performance profiles for the comparison between GMM1, GMM2, GMM3 and L-BFGS on all the 163 problems.
Figure 4: Performance profiles for the comparison between GMM1, GMM2, GMM3 and L-BFGS on the subset of problems where all the solvers find the same best function value.
Now, we consider the comparison between our methods and the state-of-the-art algorithm for unconstrained optimization, namely, L-BFGS [30]. In particular, performance profiles of solvers GMM1, GMM2, GMM3 and L-BFGS on the entire collection of 163 problems are reported in fig. 3. As we can see, our method GMM2 is quite competitive with L-BFGS in terms of computational time. Again, we have to consider the fact that the different solvers can converge toward points with different function values. In fact, as reported in table 3, GMM1 and GMM2 find a better function value more frequently than L-BFGS.
Table 3: Number of wins, i.e., the number of problems where each solver finds the best function value, on all 163 test problems. $GMM$ variants and L-BFGS solvers are considered.
Solver # wins
GMM1 55
GMM2 40
GMM3 17
L-BFGSscipy 33
In fig. 4, we thus consider the performance profiles obtained for the four solvers on the subset of problems where all of them find the same function value. As we can see, GMM2 gained some efficiency and it now slightly dominates L-BFGS in terms of computational time.
Then, for major clarity, we report in figs. 5 and 6 the performance profiles for the comparison between our bet method GMM2 and L-BFGS.
@@
Figure 5: Performance profiles for the comparison between GMM2 and L-BFGS on all the 163 problems.
Figure 6: Performance profiles for the comparison between GMM2 and L-BFGS on the subset of problems where all the solvers find the same best function value.
From figs. 5 and 6, we can say that GMM2 is quite competitive if not better than L-BFGS both in terms of computational time and number of iterations. The advantages of using GMM2 are more evident when the two solvers are compared on the subset of problems where they both find the same function value.
@@ 8 Conclusions
In this work, we introduced a general framework of gradient methods with momentum for nonconvex optimization. For the proposed class of algorithms, we proved global convergence and optimal worst-case complexity bounds. The assumptions required to obtain the theoretical results are lighter to check, from a computational perspective, than those required in related works from the literature. This result allowed to devise particularly efficient ways of implementing the proposed method. Thorough computational results we showed that the novel algorithm outperforms standard conjugate gradient methods and is competitive with L-BFGS, the state-of-the-art for nonlinear optimization problems.
Future studies might concern improvements of the proposed framework and suitable modifications to make it exploitable in different settings.
@@ Acknowledgments
The authors are very grateful to Prof. Luigi Grippo for the useful discussions at the early stages of this work.

### v3

@@ 6.2 Hessian estimation in subspace by interpolation
$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$
@@ 7 Computational experiments
In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work. We carried out the tests on two different sets of problems; the results of the two cases are presented in two different subsections.
The code for all the experiments described in this work was written in Python 3.11, exploiting numpy, scipy and pytorch libraries. The software is available at: https://github.com/gliuzzi/GMM.
Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq. 26 holds. As for the line search, we set $\delta=0.5,\;\gamma=10^{-5}$. The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration.
As a baseline for comparisons, we took into account classical nonlinear conjugate gradient methods [33] and the L-BFGS algorithm [19], the latter arguably being the most popular solver for the solution of smooth unconstrained nonlinear optimization problems. For both algorithms, we considered the very popular implementations available through the scipy library.
In a subsequent stage of our experimentation, we also took into account the cg_descent procedure available at https://people.clas.ufl.edu/hager/software/ and based on [34, 20, 33, 21]. The cg_descent provides a highly efficient implementation of a conjugate-gradient type method with subspace optimization steps. We used the python wrapper of the original C code of cg_descent v. 6.8, available at https://github.com/alexfikl/pycgdescent, so that we could employ the procedure within the same testing environment as the other considered methods.
The results of the experiments are shown in the manuscript in the form of performance profiles [35]. The detailed tables with all the result can be accessed in the code repository https://github.com/gliuzzi/GMM.
7.1 CUTEst problems
We considered a benchmark of 168 problems from the CUTEst test-suite [36]. In particular, we included in our test set all 183 problems in the CUTEst collection that are 1) unconstrained, 2) regular, 3) with a number of variables $n$ that is greater or equal than 1000 (except for the problems LUKSAN12LS, LUKSAN13LS, LUKSAN14LS having 998 variables) or that is user definable; in the latter case, we set the number of variables to 1000 or as close as possible to 1000. In Table 1 those 183 problems are reported along with their dimensions.
@@ 7.1 CUTEst problems
As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-6}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
As performance metrics, we considered both the runtime and the number of iterations. In the following, we first report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6. More in detail, we denote by GMM1, GMM2, GMM3 the algorithm where $H_{k}$ is computed as described in subsection 6.1, 6.2 and 6.3, respectively. We also included in this first test scipy conjugate gradient implementation, denoted by CGscipy, as a preliminary baseline.
Figure 1: Performance profiles on all 168 problems.
@@
In fig. 1, we report the performance profiles in terms of time and number of iterations for this comparison comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 168 problems. As we can see, the best performing solver is GMM2 both in terms of time and, together with GMM1 also in terms of iterations. Both solvers GMM2 and GMM3 are also better than CGscipy in terms of time.
@@
where $\cal S$ is the set of solvers. From fig. 2, we note that the performance of GMM methods continues to be clearly superior to that of the baseline conjugate gradient method in terms of computational time, especially in the case of GMM2.
Based on the above results, we will be considering version GMM2 in the rest of the Section as the choice for our GMM algorithm. In what follows, we also allowed a slight degree of non-monotoniticy in GMM, based on the rule proposed in [37].
We then turned to the comparison between our method and some state-of-the-art solvers for unconstrained optimization, namely, L-BFGS [19] and the cg_descent procedure [34, 20, 21]. Performance profiles of solvers GMM and L-BFGS on the entire collection of 168 problems are reported in fig. 3. As we can see, GMM seems to be slightly superior to L-BFGS in both metrics.
Figure 3: Performance profiles for the comparison between GMM and L-BFGS on all the 168 problems.
The picture heavily changes if cg_descent is included in the comparative analysis: as we can see from fig. 5, it appears by far superior than the other solvers considered in the experimentation on this benchmark of problems.
From the one hand, the superior depth and quality of implementation of the cg_descent procedure w.r.t. the simple python code implementing the GMM is probably a component that makes the gap so particularly pronounced; on the other hand, although clearly outperformed, GMM still retains a certain degree of comparability with the state-of-the-art method. This conclusion made us interested in carrying out a further comparison on a different and relevant class of problems.
@@
Figure 4: Performance profiles for the comparison between GMM, cg-descent and L-BFGS on all the 168 problems.
7.2 Multi-layer perceptron training problems
As an additional test for the proposed GMM method we considered a set of problems from supervised learning context, where methods with momentum are particularly popular. In particular, we aim to train multi-layer perceptrons for binary classification tasks [22, 23]. We considered a fully connected networks consisting of two hidden layers of 128 units each. The input layer matches the size of the samples of each dataset. The output layer consists in a single unit with sigmoid activation. The activation function of the hidden units is the hyperbolic tangent.
In total, 30 binary classification tasks are considered. We limit the dimension of the training set by taking at most 10000 samples from each dataset. We do so by randomly drawing the samples from the full dataset, as this allows to maintain the class imbalance of some datasets. A recap of the problems considered is available in table 2.
Table 2: Details of the 30 binary classification problems considered. Value $n$ denotes the number of parameters for the problem.
Dataset $n$ Dataset $n$
a1a 32001 gisette 656769
a2a 32001 ijcnn1 19585
a3a 32385 liver.disorders 17409
a4a 32385 phishing 25473
a5a 32385 sonar 24449
a6a 32385 svmguide1 17281
a7a 32513 svmguide3 19585
a8a 32513 w1a 55169
a9a 32513 w2a 55169
breast.cancer 18049 w3a 55169
cod.rna 17793 w4a 55169
covtype 23681 w5a 55169
diabetes 17793 w6a 55169
fourclass 17025 w7a 55169
german.numer 19841 w8a 55169
The weights in the multi-layer perceptron are initialized using Kaiming uniform initialization [38]. We then ran both GMM and cg_descent from the same starting point, repeating this process with three different random seeds. The objective function is binary cross-entropy. We also added an $\ell_{2}$ regularization term, with a weight $\lambda_{\textbf{reg}}=0.001$. The termination condition is $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$, which is a typical accuracy degree for learning tasks.
We report in fig. 5 the performance profiles in terms of time, number of iterations and value of the objective function obtained for GMM and cg_descent. We considered each pair dataset-seed as a distinct problem in the profiles, so that we have 90 instances overall.
@@ 7.2 Multi-layer perceptron training problems
Figure 5: Performance profiles for the comparison between GMM and cg-descent on the multi-layer perceptron training problems.
We can see from these plots that cg_descent is once again superior in terms of runtime; however, we do not observe this time differences in the number of instances solved; moreover, in a quarter of the instances GMM finished faster, and the overall performance gap appears more moderate. Actually, if we look at the number of iterations the two solver really seem comparable - the difference in runtime is arguably due to the higher number of function evaluations required by our GMM.
Another element of interest in these results concerns the final objective function value attained by the two solvers; GMM seems to have a slight advantage in this regard; we find this aspect intriguing, as as the momentum term is often used in machine learning also to escape bad stationary points in the highly nonconvex landscapes resulting from training problems.
@@ 8 Conclusions
In this work, we introduced a general framework of gradient methods with momentum for nonconvex optimization. For the proposed class of algorithms, we proved global convergence and optimal worst-case complexity bound. The assumptions required to obtain the theoretical results are lighter to check, from a computational perspective, than those required in related works from the literature. This result allowed to devise particularly efficient ways of implementing the proposed method. By thorough computational results we showed that a rather simple implementation of the novel algorithm outperforms simpler versions of conjugate gradient methods, is competitive with L-BFGS, and is somewhat comparable to state-of-the-art solvers for nonlinear optimization problems, especially in particular classes of problems, such as multi-layer perceptron training tasks.
These positive preliminary feedback on our procedure motivates future studies aimed at the further improvement of the performance of our procedure. In particular, the integration of some limited-memory and subspace optimization subroutines in cg_descent within the GMM framework might be subject of future studies.
@@ Acknowledgments
The authors are very grateful to Prof. Luigi Grippo for the useful discussions at the early stages of this work. We are also extremely thankful to the Editor-in-Chief and the anonymous referees whose very stimulating comments allowed us to improve the soundness of our work.